A conference hall, the number of seats in each row is 10 fewer than the number of rows. The number of seats in the hall are 600. Find the number of seats in each row.

Grade:11

1 Answers

Rekha Vinaya G
6 years ago
let no. of seats = x, no. of rows=y
given x=y-10  => y=x+10
total no. of seats = xy = x(x+10)
x2+10x=600
on solving, x=20